OpenStudy (anonymous):

When solving the system by the elimination method, what occurs when the second row is multiplied by 6 and the two equations are added? {4x+6y=-2 and 3x-y=9

OpenStudy (kropot72):

There is no term in y remaining. y terms have been eliminated.

OpenStudy (anonymous):

3x-y=9 Multiply by 6: 6 * ( 3x-y=9 )

OpenStudy (anonymous):

Can you multiply?

OpenStudy (kropot72):

4x + 6y = -2 18x -6y = 54 Adding gives: 22x = 52 \[x=\frac{52}{22}=2\frac{4}{11}\]